A profit-maximizing monopoly produces a good with constant marginal cost, MC = 20, that it sells in two countries. The inverse linear demand curve is P₁ = 60 Q₁ in Country 1 and P2 60 - 20₂ in = Country 2. What is the equilibrium price and quantity in each country if resale between the countries is not possible? Does the monopoly price discriminate?

1
Expert's answer
2021-12-05T19:34:47-0500

Task #274096


solution:


"MC =20"

To calculate the marginal revenue function:

Country 1 "MR =2(60Q\u2081)= 120Q\u2081"


Country 2 "MR= 60-2(2Q\u2082) =60-4Q\u2082"


To find the equilibrium quantity in each country, equate:


"MR=MC"


For country 1

"120Q\u2081 =20"

"Q=0.167"


Equilibrium price will be found by substituting Q in the inverse demand function.


"P1=60 (0.167)"

"P1=10"


Country 2

"MR =MC"

"60-4Q\u2082=20"

"Q\u2082=10"


Equilibrium price will be found by substituting Q in the inverse demand function.

"P2=60 - 2(10)=40"

"P2=40"


Yes there is price discrimination because, the same vendor sells identical goods or services at different prices in the market. The seller's purpose in price discrimination is to make the most money feasible. Despite the fact that the cost of manufacturing the products is the same, the seller chose to raise the price based on geography, consumer financial status, and product demand.